OpenStudy (anonymous):

Triangle XYZ has vertices X(1,2), Y(3,6), Z(-1,4) Use a determinant to find the area of XYZ

OpenStudy (anonymous):



OpenStudy (anonymous):

So you get 1(6-4) -2(3+1) +1(12+6)=2-6+18=14

OpenStudy (anonymous):

@shankvee you should put ABS before the determinant